# Simple Controller Design to Achieve Iso-Damping Robustness: Non-Iterative Data-Driven Approach Based on Fractional-Order Reference Model

### v2

@@ Abstract
This study proposes a simple controller design approach to achieve a class of robustness, the so-called iso-damping property. The proposed approach can be executed using only one-shot input/output data. An accurate mathematical model of a controlled plant is not required. The model-reference control problem is defined to achieve the desired closed-loop specifications, including the iso-damping, and the reference model is designed on the basis of fractional-order calculus. The optimization problem for the model-reference control is formulated using the one-shot input/output data while considering the bounded-input bounded-output (BIBO) stability from a bounded reference input to a bounded output. The iso-damping robust controller is obtained by solving the optimization problem. The representative advantages of the proposed approach over the conventional methods are the simplicity, practicality, and reliability from the viewpoint of the unnecessity of the plant model and explicit consideration of the BIBO stability from a bounded reference input to a bounded output. Numerical examples demonstrate the validity of the proposed approach.
@@ I-A Motivation
Robustness is a fundamental requirement for control systems in practice, since the controlled plant has various uncertainties and characteristic changes. Numerous studies have been conducted to achieve robust control [1]. A conventional approach to ensure robustness is 1) modeling the nominal plant and assuming the range of the uncertainty and 2) designing the controller based on the model-based control theory to work well in the pre-specified uncertainty range. However, this approach fails if the plant model including uncertainty is not appropriate [2]; modeling the nominal plant and the uncertainty relies on the expert’s knowledge, increasing the development cost. Therefore, establishing a practical, simple, and reliable framework for robust controller design is an important challenge in the control engineering community.
Among the various types of uncertainties, gain variation is often found in practical control systems (e.g., load changes in robotic systems [3]). Gain variation adversely affects the control performance and even causes instability. Thus, the robustness to plant gain variation is an indispensable requirement for practical control systems. Recently, the iso-damping property [4] has garnered a great deal of attention. The iso-damping property is a special class of robustness: the invariance of the overshoot of the setpoint response under plant gain variation. Hence, the iso-damping property is highly desirable for practical control systems, enhancing the reliability and safety.
@@
If the Bode phase plot of the open-loop transfer function is locally flat around the gain crossover frequency, such a special characteristic is called the flat-phase property [4]. A feedback control system with the flat-phase property achieves the iso-damping robustness [4]. $\blacksquare$
I-B Related works
@@
Control systems involving a fractional-order calculus (FC) has generated a great deal of attention: fractional-order (FO) system theory (e.g., stabilization of FO systems [5, 6, 7, 8, 9], stability tests [10, 11, 12, 13], the generalized KYP lemma for singular FO systems [14], the FC-based gradient algorithm [15]). FC is a mathematical framework generalizing traditional integer-order (IO) calculus, which considers non-integer order differentiations and integrations. Focusing on its rich expression ability and generality, Magin et al. have discussed the benefit of exploiting the idea of FC in the cybernetics field [16]. In particular, FO control is a powerful and effective approach to achieving good control performance and robustness [17]. For example, the Bode’s ideal transfer function (BITF) is an FO transfer function with the ideal flat-phase specification; it is often utilized for controller design to achieve the iso-damping robustness [18]. FC-based robust control is one important direction of the exploitation of FC in the cybernetics field.
@@
Various approaches have been proposed to achieve iso-damping on the basis of FO control. These methods aim to tune some design parameters of the fixed structure controller, such as the FO proportional-integral-derivative (PID) controller, so as to realize the flat-phase property. The analytical conditions for the flat-phase have been derived for a special class of plants and controllers: the FO-PID controller for simplified FO plants with time delay [19], the FO-PID controller for stable first-order plus time delay (FOPTD) plant [20], and the FO-PID controller with FO lowpass filter [21]. The graphical approach (e.g., plotting the feasible parameter region or achievable performance specification) is an effective way to determine controller parameters satisfying the analytical flat-phase condition. As for the graphical approach, one study has utilized the Bode phase plot [22], one has proposed a flat-phase curve [23], and a novel procedure to tune the FO-[PID] controller has been established for the FOPTD plant [24]. It should be noted that both the analytical and graphical approaches require accurate mathematical models and parameters of the actual controlled plant. However, the necessity of the plant model and parameters imposes heavy burdens on designers and causes the technique to be fail under modeling errors. Moreover, in most cases, the analytical and graphical policies are derived for very simple plants such as the FOPTD system. Thus, not only the plant model but also its bold approximation is necessary to employ such a method. Controller design based on the approximated model often causes severe performance degradation due to the gap between the actual plant and the reduced one for controller design.
@@
Numerical optimization-based approaches have been examined for a wide variety of plants in the context of parameter tuning of FO control systems (e.g., the differential evolution algorithm [25], the wolf based elephant herding method [26]). As for realizing the flat-phase specification, a novel metaheuristics algorithm has been developed for FO-PID controller tuning [27]. Nevertheless, the conventional optimization-based design requires an accurate plant model and parameters for evaluating the objective function.
@@
Recently, controller design techniques based on the data rather than the mathematical model of the plant have attracted a great deal of attention: direct data-driven control (D3C) [28, 29, 30]. Exploiting the D3C approach to design an FO control system can overcome the drawbacks of the abovementioned challenges, since these problems are mainly raised by the needs for accurate plant models and parameters. Several studies have employed frequency response data for designing FO controllers achieving the iso-damping property [31, 32]. However, obtaining the frequency response is cumbersome and often requires a special experimental setup.
@@
On the other hand, to the best of our knowledge, few studies have explored an MR-D3C approach for FO control. The IFT technique has been employed to tune the FO-PID controller for a hydraulic actuator [44]. However, the tuning scheme presented in [44] requires multiple experiments, imposing heavy burdens on designers. The FO-PID controller has been tuned on the basis of the VRFT [45, 46, 47]. Nevertheless, the conventional VRFT does not consider closed-loop stability. Although one study has tuned the FO-PID controller based on the MR-D3C approach, it does not consider the iso-damping robustness [48]. Consequently , to the best of our knowledge, no studies have proposed simple and practical FO MR-D3C techniques especially for achieving the iso-damping property.
@@ I-C Contribution and novelty
This study presents a novel controller design scheme for a discrete-time single-input single-output (SISO) linear time-invariant (LTI) system. The proposed approach does not require a mathematical model of the controlled plant. The present approach can be executed using only one-shot input/output data collected by a closed-loop experiment. The resultant controller due to the proposed approach can achieve iso-damping: the overshoot amount of the closed-loop system is almost invariant under the plant gain variation. In the present approach, the robust controller design problem is defined as the MR-D3C problem using the reference model a priori chosen by the designer. The reference model is designed on the basis of an FO system, defining the desired characteristics of the closed-loop system such as the flat-phase property. The MR-D3C problem is reformulated as an optimization problem based on a fictitious reference signal (the fictitious reference signal is computed using the one-shot input/output data). The loss function to be minimized reflects the information of the closed-loop poles from a reference input to an output. The solution of the minimization problem yields a controller achieving the desired closed-loop characteristic such as the iso-damping specified by the reference model. Numerical examples demonstrate the validity of the proposed approach.
Compared with conventional approaches focusing on iso-damping robust control, the contribution and novelty can be summarized as follows:
@@
    (Simplicity) The proposed approach does not require the mathematical model of the actual controlled plant, providing freedom from costly and burdensome modeling procedures. Therefore, the present approach is simpler than model-based approaches such as analytical and graphical approaches [19, 20, 21, 22, 23, 24] and numerical optimization based on the plant model [18, 27, 25].
@@
    (Reduced burdens on designers) The appropriate controller is automatically obtained via solving the optimization problem in the proposed approach. The manual trial-and-error requiring the expert’s skill is unnecessary. Moreover, the proposed tuning scheme is executed on the basis of only one-shot input/output data, unlike the traditional FO MR-D3C technique requiring multiple datasets [44]. Thus, the proposed approach significantly reduces the burdens on the designer.
@@
    (Reliability) Unlike the conventional FO MR-D3C techniques [45, 46, 47], information regarding the closed-loop pole is explicitly reflected in the proposed optimization problem for the controller design. This feature allows the present technique to explicitly consider the bounded-input bounded-output (BIBO) stability of the closed-loop system from a bounded reference input to a bounded output. Therefore, the proposed design approach yields a reliable controller.
@@
    (Generality) Most analytical approaches are developed for a special class of plants and controllers, as discussed above. On the other hand, the proposed approach is applicable to a wide variety of discrete-time SISO LTI plants and controllers.
@@ I-D Structure of this paper
This paper is organized as follows. Section II summarizes some preliminaries and states the problem addressed in this study. Section III describes the proposed controller design approach. We show the numerical studies to demonstrate the validity of the proposed approach in Section IV. The discussion is given in Section V. Section VI concludes the paper.
@@ II Preliminaries
$\mathrm{Tpl}(\it{v})\coloneqq\begin{bmatrix}v_{1}&0&0&\ddots&0\\ \vdots&v_{1}&0&\ddots&0\\ v_{n-2}&\ddots&v_{1}&0&\vdots\\ v_{n-1}&v_{n-2}&\ddots&\ddots&0\\ v_{n}&v_{n-1}&v_{n-2}&\ldots&v_{1}\\ \end{bmatrix}.$ (1)
Here, $v_{j}\;(j=1,2,\ldots n)$ is the $j$-th component of $v$. For an $n$-length sequence $v=\{v_{k}\}_{k=0}^{n-1}\;(1\leq n\leq\infty)$, $\mathrm{vec}(v)\coloneqq\begin{bmatrix}v_{0}&v_{1}&\ldots&v_{n-1}\end{bmatrix}^{\top}$ and $\mathrm{Tpl}\circ\mathrm{vec}\coloneqq\mathrm{Tpl}\left(\mathrm{vec}(\cdot)\right)$. For a matrix $M\in\mathbb{R}^{n\times n}$, $\Lambda\left(M\right)$ denotes the set of all eigenvalues of $M$ including their algebraic multiplicity. The spectral radius of $M$ is represented as $\rho\left(M\right)$, i.e., $\rho(M)\coloneqq\max_{\lambda\in\Lambda(M)}\left\lvert\lambda\right\rvert$.
[…]
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI system and let $\{g_{k}\}_{k=0}^{\infty}$ be the impulse response sequence of $G\left(z\right)$. Then, the output $y=\{y_{k}\}_{k=0}^{\infty}$ of $G\left(z\right)$ to the input $u=\{u_{k}\}_{k=0}^{\infty}$ is $y_{t}=g_{t}\ast u_{t}\;(t=0,1,2,\ldots)$. Here, the symbol $\ast$ denotes the convolution, i.e., $g_{t}\ast u_{t}\coloneqq\sum_{n=0}^{t}g_{\tau}u_{t-\tau}=\sum_{n=0}^{t}g_{t-\tau}u_{\tau}$ [49]. We sometimes denote the input/output relation as $y=G(z)u$.
The $p$-norm $\left\lVert v\right\rVert_{p}$ of $v\in\mathbb{R}^{n}$ is defined as $\left\lVert v\right\rVert_{p}\coloneqq\left(\sum_{k=1}^{n}\left\lvert v_{j}\right\rvert^{p}\right)^{\frac{1}{p}}$ for $1\leq p<\infty$ and $\left\lVert v\right\rVert_{\infty}\coloneqq\sup_{j}\left\lvert v_{j}\right\rvert$. Similarly, the $l_{p}$-norm $\left\lVert u\right\rVert_{p}$ of a real-valued signal $u=\{u_{k}\}_{k=0}^{\infty}$ is defined as $\left\lVert u\right\rVert_{p}\coloneqq\left(\sum_{k=1}^{\infty}\left\lvert u_{j}\right\rvert^{p}\right)^{\frac{1}{p}}$ for $1\leq p<\infty$ and $\left\lVert u\right\rVert_{\infty}\coloneqq\sup_{j}\left\lvert u_{j}\right\rvert$.
II-A The flat-phase property and iso-damping

### v3

@@ Abstract
This study proposes a simple controller design approach to achieve a class of robustness, the so-called iso-damping property. The proposed approach can be executed using only one-shot input/output data. An accurate mathematical model of a controlled plant is not required. The model-reference control problem is defined to achieve the desired closed-loop specifications, including the iso-damping, and the reference model is designed on the basis of fractional-order calculus. The optimization problem for the model-reference control is formulated using the one-shot input/output data while considering the bounded-input bounded-output (BIBO) stability from a bounded reference input to a bounded output. The iso-damping robust controller is obtained by solving the optimization problem. The representative advantages of the proposed approach over the conventional methods are the simplicity, practicality, and reliability from the viewpoint of the unnecessity of the plant model and explicit consideration of the BIBO stability from a bounded reference input to a bounded output. Numerical and experimental studies demonstrate the validity of the proposed approach.
@@ I-A Motivation
Robustness is a fundamental requirement for control systems, as the controlled plant has various uncertainties and characteristic changes. Numerous studies have been conducted to achieve robust control [1]. A conventional approach to ensure robustness is 1) modeling the nominal plant and assuming the range of the uncertainty and 2) designing the controller based on the model-based control theory to work well in the pre-specified uncertainty range. However, this approach fails if the plant model including uncertainty is not appropriate [2]; modeling the nominal plant and the uncertainty relies on the expert’s knowledge, increasing the development cost. Therefore, establishing a practical, simple, and reliable framework for robust controller design is an important challenge.
Among the various uncertainties, gain variation is often found in practical control systems (e.g., load changes in robotic systems [3]). Gain variation adversely affects the control performance and even causes instability. Thus, the robustness to plant gain variation is an indispensable requirement in practical applications. Recently, the iso-damping property [4] has garnered a great deal of attention. The iso-damping property is a special class of robustness: the invariance of the overshoot of the setpoint response under plant gain variation. Hence, the iso-damping property is highly desirable for practical control systems, enhancing the reliability and safety.
@@
If the Bode phase plot of the open-loop transfer function is locally flat around the gain crossover frequency, such a special characteristic is called the flat-phase property [4]. A feedback control system with the flat-phase property achieves the iso-damping robustness [4].
I-B Related work
@@
Control systems involving a fractional-order calculus (FC) has generated a great deal of attention: fractional-order (FO) system theory (e.g., stabilization of FO systems [5, 6, 7, 8, 9], stability tests [10, 11, 12, 13], the generalized KYP lemma for singular FO systems [14], the FC-based gradient algorithm [15]). FC is a mathematical framework generalizing traditional integer-order (IO) calculus, which considers non-integer order differentiations and integrations. Focusing on its rich expression ability and generality, Magin et al. have discussed the benefit of exploiting the idea of FC in the cybernetics field [16]. In particular, FO control is a powerful and effective approach to achieving good control performance and robustness [17]. The advantages of FO control have been demonstrated in various industrial applications (e.g., medical devices [18], thermal power systems [19]), highlighting its practical value. The Bode’s ideal transfer function (BITF) is an FO transfer function with the ideal flat-phase specification; it is often utilized for controller design to achieve the iso-damping robustness [20]. FC-based robust control is one important direction of the exploitation of FC in the cybernetics field.
@@
Various approaches have been proposed to achieve iso-damping on the basis of FO control. These methods aim to tune some design parameters of the fixed structure controller, such as the FO proportional-integral-derivative (PID) controller, in orders to realize the flat-phase property. The analytical conditions for the flat-phase have been derived for a special class of plants and controllers [21, 22, 23]. The graphical approach (e.g., plotting the feasible parameter region or achievable performance specification) is an effective way to determine controller parameters satisfying the analytical flat-phase condition [24, 25, 26]. These approaches have been established by the frequency domain analyses. It should be noted that both the analytical and graphical approaches in the frequency-domain require accurate mathematical models and parameters of the actual controlled plant. However, the necessity of the plant model and parameters imposes heavy burdens on designers; modeling errors can lead to unexpected deterioration of control performance. Moreover, in most cases, the analytical and graphical policies are derived for very simple plants such as the first-order plus time delay system. Thus, not only the plant model but also its bold approximation is necessary to employ such a method. Controller design based on the approximated model often causes performance degradation due to the gap between the actual plant and the reduced one for controller design.
@@
Numerical optimization-based approaches have been examined for a wide variety of plants in the context of parameter tuning of FO control systems [27, 28]. As for realizing the flat-phase specification, a novel metaheuristic algorithm has been developed for FO-PID controller tuning [29]. The process of the conventional optimization-based controller tuning can be summarized as follows: (1) perform closed-loop control tests; (2) assess the performance criteria using the results from step (1); (3) determine the next candidate controller based on the evaluation results and the optimization approach; and (4) iterate steps (1)–(3). Nevertheless, this procedure implies that the conventional optimization-based tuning requires huge amount of closed-loop control tests. Conducting control experiments in a real system is time-consuming; simulation-based closed-loop tests require an accurate plant model and parameters for evaluating the objective function.
@@
Recently, controller design techniques based on the data rather than the mathematical model of the plant have attracted a great deal of attention: direct data-driven control (D3C) [30, 31, 32]. Exploiting the D3C approach to design an FO control system can overcome the drawbacks of the abovementioned challenges, since these problems are mainly raised by the needs for accurate plant models and parameters. Several studies have employed frequency response data for designing FO controllers [33, 34]. However, obtaining the frequency response data is cumbersome and often requires a special experimental setup.
@@
On the other hand, to the best of our knowledge, few studies have explored an MR-D3C approach for FO control. The IFT technique has been employed to tune the FO-PID controller for a hydraulic actuator [46]. However, the tuning scheme presented in [46] requires multiple experiments, imposing heavy burdens on designers. The FO-PID controller has been tuned on the basis of the VRFT [47, 48, 49]. Nevertheless, the conventional VRFT does not consider closed-loop stability. One study [50] has proposed a VRFT-based FO controller tuning method with stability considerations; however, the stability verification requires the frequency response data of the controlled plant, which increases the user’s burden. Although one study has tuned the FO-PID controller based on the MR-D3C approach, it does not consider the iso-damping robustness [51]. Consequently, to the best of our knowledge, no studies have proposed simple and practical FO MR-D3C techniques especially for achieving the iso-damping property.
@@ I-C Contribution and novelty
[h]
TABLE I: Comparison between the proposed approach and the conventional approaches for iso-damping robustness based on FO system.
Advantage
Challenge
Conventional      Model-based synthesis in frequency-domain Computationally inexpensive Necessity of plant model Applicable to simple plants and controllers only         Frequency response data-based tuning Free from the necessity of plant model Need for special experimental setup to obtain frequency response data         Minimization of time- domain performance index Broad applicability Free from the necessity of plant model $\dagger$} Need for iterative closed-loop control tests Involving non-convex optimization    Proposed      Data-driven approach using fictitious reference Free from the necessity of plant model Broad applicability Easy data collection $\ddagger$} Involving non-convex optimization
  • $\dagger$}
    This advantage is realized when the performance index is evaluated based on real-world experiments.
  • $\ddagger$}
    The proposed approach requires only a single set of time-series input-output data from the controlled plant.
This study presents a novel controller design scheme for a discrete-time single-input single-output (SISO) linear time-invariant (LTI) system. The proposed approach does not require a mathematical model of the controlled plant. The present approach can be executed using only one-shot input/output data. The resultant controller due to the proposed approach can achieve iso-damping: the overshoot amount of the closed-loop system is almost invariant under the plant gain variation. In the present approach, the robust controller design problem is defined as the MR-D3C problem using the reference model a priori chosen by the designer. The reference model is designed on the basis of an FO system, defining the desired characteristics of the closed-loop system such as the flat-phase property. The MR-D3C problem is reformulated as an optimization problem based on a fictitious reference signal (the fictitious reference signal is computed using the one-shot input/output data). The loss function to be minimized reflects the information of the closed-loop poles from a reference input to an output. The solution of the minimization problem yields a controller achieving the desired closed-loop characteristic such as the iso-damping specified by the reference model. The validity of the proposed approach is demonstrated through numerical and experimental studies.
Table I-C highlights the distinctions and advancements of the proposed approach over the conventional approaches for iso-damping robust control using FO systems. Compared with the conventional approaches, the contribution and novelty of this study can be summarized as follows:
@@
    (Simplicity) The proposed approach does not require the mathematical model of the actual controlled plant, providing freedom from costly and burdensome modeling procedures. Therefore, the present approach is simpler than model-based approaches such as analytical and graphical approaches [21, 22, 23, 24, 25, 26] and numerical optimization based on the plant model [20, 29, 27]. Moreover, the proposed approach is based not on frequency response data but on time-series data, which makes the proposed approach more straightforward than the frequency response data-based approaches [34, 33]. This is because obtaining frequency response data is more difficult than obtaining time-series data, as it often requires a special experimental setup.
@@
    (Reduced burdens on designers) The appropriate controller is automatically obtained via solving the optimization problem in the proposed approach. The manual trial-and-error requiring the expert’s skill is unnecessary. Moreover, the proposed tuning scheme is executed on the basis of only one-shot input/output data, unlike the traditional FO MR-D3C technique requiring multiple closed-loop tests such as the IFT-based approach [46]. Thus, the proposed approach significantly reduces the burdens on the designer.
@@
    (Reliability) Unlike the conventional FO MR-D3C techniques based on the VRFT [47, 48, 49], information regarding the closed-loop pole is explicitly reflected in the proposed optimization problem for the controller design. This feature allows the present technique to explicitly consider the bounded-input bounded-output (BIBO) stability of the closed-loop system from a bounded reference input to a bounded output. Therefore, the proposed design approach yields a reliable controller.
@@
    (Generality) Most analytical approaches and frequency response data-based techniques are developed for a special class of plants and controllers, as discussed above. On the other hand, the proposed approach is applicable to a wide variety of discrete-time SISO LTI plants and controllers.
@@ I-D Structure of this paper
This paper is organized as follows. Section II summarizes some preliminaries and states the problem addressed in this study. Section III describes the proposed controller design approach. We show the numerical and experimental studies to demonstrate the validity of the proposed approach in Section IV. The discussion is given in Section V. Section VI concludes the paper.
@@ II Preliminaries
$\mathrm{Tpl}(\it{v})\coloneqq\begin{bmatrix}v_{1}&0&0&\ddots&0\\ \vdots&v_{1}&0&\ddots&0\\ v_{n-2}&\ddots&v_{1}&0&\vdots\\ v_{n-1}&v_{n-2}&\ddots&\ddots&0\\ v_{n}&v_{n-1}&v_{n-2}&\ldots&v_{1}\\ \end{bmatrix}$ (1)
where $v_{j}\;(j=1,2,\ldots n)$ is the $j$-th component of $v$. For an $n$-length sequence $v=\{v_{k}\}_{k=0}^{n-1}\;(1\leq n\leq\infty)$, $\mathrm{vec}(v)\coloneqq\begin{bmatrix}v_{0}&v_{1}&\ldots&v_{n-1}\end{bmatrix}^{\top}$ and $\mathrm{Tpl}\circ\mathrm{vec}\coloneqq\mathrm{Tpl}\left(\mathrm{vec}(\cdot)\right)$. For a matrix $M\in\mathbb{R}^{n\times n}$, $\Lambda\left(M\right)$ denotes the set of all eigenvalues of $M$ including their algebraic multiplicity. The spectral radius of $M$ is denoted as $\rho\left(M\right)$, i.e., $\rho(M)=\max_{\lambda\in\Lambda(M)}\left\lvert\lambda\right\rvert$.
The Laplace variable and the $Z$-variable are represented as $s$ and $z$, respectively. If a transfer function $G\left(\cdot\right)$ has the tunable parameter $\theta\in\mathbb{R}^{n}$, we sometimes denote the transfer function as $G\left(\cdot;\theta\right)$. For a continuous-time transfer function $G\left(s\right)$, $\mathrm{c2d}\left(G\left(s\right)\right)$ represents its discretization. We write $\mathrm{f2i}\left(G_{FO}\left(s\right)\right)$ as the IO approximation of the FO transfer function $G_{FO}\left(s\right)$. Also, $\mathrm{c2d}\circ\mathrm{f2i}\left(\cdot\right)\coloneqq\mathrm{c2d}\left(\mathrm{f2i}\left(\cdot\right)\right)$ represents the transformation from a continuous-time FO transfer function to the corresponding discrete-time IO transfer function.
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI system and let $\{g_{k}\}_{k=0}^{\infty}$ be the impulse response sequence of $G\left(z\right)$. Then, the output $y=\{y_{k}\}_{k=0}^{\infty}$ of $G\left(z\right)$ to the input $u=\{u_{k}\}_{k=0}^{\infty}$ is $y_{t}=g_{t}\ast u_{t}\;(t=0,1,2,\ldots)$. Here, the symbol $\ast$ denotes the convolution, i.e., $g_{t}\ast u_{t}\coloneqq\sum_{\tau=0}^{t}g_{\tau}u_{t-\tau}=\sum_{\tau=0}^{t}g_{t-\tau}u_{\tau}$ [52]. We sometimes denote the input/output relation as $y=G(z)u$.
The $p$-norm $\left\lVert v\right\rVert_{p}$ of $v\in\mathbb{R}^{n}$ is defined as $\left\lVert v\right\rVert_{p}\coloneqq\left(\sum_{k=1}^{n}\left\lvert v_{j}\right\rvert^{p}\right)^{\frac{1}{p}}$ for $1\leq p<\infty$ and $\left\lVert v\right\rVert_{\infty}\coloneqq\sup_{j}\left\lvert v_{j}\right\rvert$. Similarly, the $l_{p}$-norm $\left\lVert u\right\rVert_{p}$ of a real-valued signal $u=\{u_{k}\}_{k=0}^{\infty}$ is defined as $\left\lVert u\right\rVert_{p}\coloneqq\left(\sum_{k=0}^{\infty}\left\lvert u_{j}\right\rvert^{p}\right)^{\frac{1}{p}}$ for $1\leq p<\infty$ and $\left\lVert u\right\rVert_{\infty}\coloneqq\sup_{j}\left\lvert u_{j}\right\rvert$.
II-A Flat-phase property and iso-damping